# Open Agent Specification (Agent Spec) Technical Report

### v3

@@ Abstract
Open Agent Specification (Agent Spec) is a declarative language that allows AI agents and their workflows to be defined in a way that is compatible across different AI frameworks, promoting portability and interoperability within AI Agent frameworks.
Agent Spec aims to resolve the challenges of fragmented agent development by providing a common unified specification that allows AI agents to be designed once and deployed across various frameworks, improving interoperability and reusability, and reducing redundant development efforts. Additionally, Agent Spec facilitates development tools and portability, allowing AI agents to be defined independently of their execution environment and enabling teams to exchange solutions without implementation-specific limitations.
Agent Spec benefits four key groups: (i) Agent developers, who gain access to a superset of reusable components and design patterns, enabling them to leverage a broader range of functionalities; (ii) Agent framework and tool developers, who can use Agent Spec as an interchange format and therefore benefit from the support of other frameworks as well as other tools; (iii) Researchers, who can achieve reproducible results and comparability, facilitating more reliable and consistent outcomes; (iv) Enterprises, which benefit from faster prototype-to-deployment, increased productivity, as well as greater scalability and maintainability for their AI agent solutions. This technical report provides an overview of the technical foundations of Agent Spec, including motivation, benefits, and future developments.
Agent Spec also introduces a standardized Evaluation harness to assess agent behavior and agentic workflows across runtimes - analogous to how HELM and related harnesses standardized LLM evaluation - so that performance, robustness, and efficiency can be compared consistently across frameworks. We demonstrate this using four distinct runtimes (LangGraph, CrewAI, AutoGen, and WayFlow) evaluated over three different benchmarks (SimpleQA Verified, $\tau^{2}$-Bench and BIRD-SQL).
@@ 1 Introduction
This technical report introduces Agent Spec, a declarative, framework-agnostic configuration language designed to define AI agents and their workflows with high fidelity. By providing a unified representation, Agent Spec enables seamless portability and interoperability of agents across diverse frameworks, ensuring consistent behavior and execution across various AI Agent frameworks. Agent Spec serves as an abstraction layer above framework-specific specifications, acting as a unifying layer that encapsulates agent functionality beyond individual framework constraints. It also serves as an Evaluation harness, enabling consistent comparison of agent and flow patterns across different frameworks (e.g., LangGraph, CrewAI, AutoGen, and WayFlow). Using this capability, we report initial results on representative tasks from three benchmark datasets (SimpleQA Verified [4], $\tau^{2}$-Bench [2] and BIRD-SQL [14]).
1.1 Background & Motivation
@@
Existing agentic frameworks have their own strengths and limitations. While flexibility is crucial to accommodating diverse agentic design patterns, currently available frameworks do not fully support all of them. Table 1 outlines key objectives a comprehensive framework should cover and describes how Agent Spec addresses each of these requirements.
@@
Table 1: Challenges and objectives in agentic frameworks and solutions that we define with Agent Spec.
@@
Agents
Author and edit conversational solutions
such as ReAct [15] style agents.
Ability to define components that represent "resources"
(with no impact on conversation flow routing) such as
LLMs, memory, prompt templates, tools, other Agents
which can be attached to other components’ properties.
Flows
Author and edit graph-style workflows
such as multi-step business processes.
Ability to define components that have a well-defined
execution path such as a directed graph which may
contain conditional routes or loops.
Control flow routing
Ability to define the flow of execution
between components in the system,
including conditional or looping flows.
Explicit "relationships" (edges) for control flow
supporting one-to-many and many-to-one transitions. The
specific transition to be taken is determined at
runtime for a given execution.
I/O routing
Ability to define which outputs are
consumed by which inputs for fixed
flow components. This includes
conditional or looping flows, and outputs
that are consumed in multiple places.
Explicit "relationships" (edges) for data flow. Support
for one-to-many and many-to-one transitions where the specific
transition to be taken is determined at runtime for a
given execution.
Nesting/Composition
Reuse components or solutions at any
level of granularity in other solutions.
Includes multi-agent composition.
Definition of components that can be encapsulated and
reused elsewhere. Established required properties for
such components. Varying levels of visibility of
exchanged messages between components.
Orchestration of execution order and dependency either
statically or dynamically.
Referencing
Ability to define a component instance
once and refer to it in multiple places.
"Reference" syntax to specify property values by
referencing and reuse. Provision of components for
defining an element and for using them.
For example, an Agent component to define agent and
an AgentNode in a Flow to use it in a certain place,
or a Tool component to define a tool and a ToolNode
in a Flow to use the tool in a specific place.
@@
Moreover, different frameworks are usually parameterized and configured by different means. This makes porting solutions between frameworks a tedious and error-prone process, hindering collaboration and knowledge sharing among teams developing agentic applications. Consequently, a declarative, portable, and framework-agnostic representation of agentic designs is needed.
For instance, ONNX [10] has revolutionized deep learning by providing a consistent way to port ML models across different frameworks (e.g., PyTorch [11], TensorFlow [13]). Similarly, Agent Spec aims to establish a unified representation for AI agents, enabling seamless interoperability and execution across diverse agentic frameworks. Just as ONNX allows models to be trained in one framework and executed in another, Agent Spec allows AI agents to be designed once and deployed across multiple platforms without reimplementation.
@@ 1.2 Objectives of Open Agent Specification
Agent Spec aims to define a common representation and act as a superset of capabilities for AI agents and workflows. Agent Spec graphs can be executed in different frameworks by means of runtimes that map Agent Spec components to the underlying framework’s primitives. Agent Spec, along with importers and exporters for various underlying frameworks, enables transparent portability of agent workflows. Agent Spec also provides methods to validate agent workflows before execution, ensuring they are syntactically and logically correct and executable.
1.3 Key Benefits of Agent Spec
@@ 1.3.1 Portability & Reusability
  • Framework-Agnostic: Agent Spec abstracts agent definitions from specific implementations by providing a unified declarative representation, allowing agents to operate uniformly across multiple platforms (e.g., AutoGen, LangGraph, OCI Agents) without any need for reimplementation.
  • Modular Design: Agent Spec’s component-based structure fosters reusability and extensibility, enabling enterprises to build complex agentic systems by assembling and configuring standardized components.
  • Support for Complex Agentic Flows: Agent Spec facilitates the creation of intricate agentic workflows by supporting various modular components and reusable Agent Spec-defined patterns, enabling streamlined development of advanced AI agent systems.
  • Portable Execution Graphs: Agent Spec enables execution graphs to be described in a way that is independent of the underlying runtime implementation, ensuring seamless portability across frameworks. Converters between the representations of other frameworks to Agent Spec would simplify the porting of previously developed agents to Agent Spec-compatible frameworks.
@@ 1.3.2 Interoperability & Compatibility
  • Robustness & Consistency: Agents defined in Agent Spec can be built on top of components that have been defined and validated within the Agent Spec specification, leading to higher reliability and predictability.
  • Define Once, Run on Multiple Frameworks: Agentic systems leveraging Agent Spec can run on various frameworks through runtime adapters. Agent Spec’s open specification allows developers to implement these adapters to support any framework, enabling agents to run in optimized environments.
@@ 1.3.3 Unification across Frameworks
  • Common Format: Agent Spec’s common format simplifies integration of agents and enables easier agent sharing across platforms.
  • Knowledge Exchange: The unified declarative format that Agent Spec provides helps developers as well as maintainers of agents to efficiently exchange solutions and ideas.
  • Conformance Test Suite: The Agent Spec conformance test suite offers a comprehensive set of tests that cover the various functionalities and components available within the specification, such as Agent, Flow, and ToolNode. It includes illustrative validation examples of Agent Spec YAML configurations and verifies that these configurations produce consistent execution outcomes and behaviors across multiple runtime environments. This approach ensures interoperability and correctness of implementations supporting the Agent Spec specification.
2 Open Agent Specification Design
This section provides an overview of Agent Spec’s declarative model. It details how Agent Spec functions across different frameworks and outlines implementation specifics, including the SDK structure, supported programming languages, agent definition using the SDK, and the serialization/deserialization of agents and workflows.
Agent Spec is intended to be a portable, platform-agnostic configuration language that allows agents and agentic systems to be described with sufficient fidelity. It defines the conceptual objects, called components, that compose agents in typical agentic systems, including the properties that determine the components’ configuration, and their respective semantics.
Runtimes implement the Agent Spec components for execution with agentic frameworks or libraries. Agent Spec would be supported by SDKs in various languages (e.g., Python), enabling serialization and deserialization of Agents to YAML, or creating them from object representations with assurance of conformance to the specification.
An outline of the components involved in the Agent Spec ecosystem is depicted in Fig. 1.
Table 2: Agent Spec aims to do for AI agents what ONNX did for ML models
@@ 2 Open Agent Specification Design
Scope
Unified representation of
ML models, allowing portability
across different deep learning
frameworks.
Unified representation of AI agents
and workflows, enabling them to
run across diverse agentic frameworks.
Portability
ONNX allows ML models to be
trained in one framework (e.g.,
PyTorch, TensorFlow) and
executed in another.
Agent Spec enables AI agents to be built in
one framework (e.g., LangGraph [12], AutoGen [9],
OCI Agents [7]) and run in another without modification.
Unification
Provides a common declarative
format which is versioned as
an Opset.
Provides a common agent
configuration format.
Extensibility
Supports various ML
operations and optimizations
Defines modular components for scalable agent systems.
Figure 1: Agent Spec’s Design
2.1 Overview of Agent Specification
The Open Agent Specification defines structure and behavior of the conceptual building blocks, called components, that make up agents in typical agent-based systems. This structure can be serialized into a serialization language, like JSON or YAML. An overview of the language specification of Agent Spec follows, while the full specification can be accessed at https://oracle.github.io/agent-spec/agentspec/language_spec_25_4_1.html.

### v4

@@ Abstract
The proliferation of agent frameworks has led to fragmentation in how agents are defined, executed, and evaluated. Existing systems differ in their abstractions, data flow semantics, and tool integrations, making it difficult to share or reproduce workflows. We introduce Open Agent Specification (Agent Spec), a declarative language that defines AI agents and agentic workflows in a way that is compatible across frameworks, promoting reusability, portability and interoperability of AI agents. Agent Spec defines a common set of components, control and data flow semantics, and schemas that allow an agent to be defined once and executed across different runtimes. Agent Spec also introduces a standardized Evaluation harness to assess agent behavior and agentic workflows across runtimes - analogous to how HELM and related harnesses standardized LLM evaluation - so that performance, robustness, and efficiency can be compared consistently across frameworks. We demonstrate this using four distinct runtimes (LangGraph, CrewAI, AutoGen, and WayFlow) evaluated over three different benchmarks (SimpleQA Verified, $\tau^{2}$-Bench and BIRD-SQL). We provide accompanying toolsets: a Python SDK (PyAgentSpec), a reference runtime (WayFlow), and adapters for popular frameworks (e.g., LangGraph, AutoGen, CrewAI). Agent Spec bridges the gap between model-centric and agent-centric standardization & evaluation, laying the groundwork for reliable, reusable, and portable agentic systems.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
The rapid emergence of AI agent frameworks has accelerated the development of systems capable of planning, invoking tools, and coordinating multi-step workflows [SAPKOTA2026103599; orclblog; xie2025ai; wolflein-etal-2025-llm]. However, this progress has also introduced fragmentation: each framework defines its own abstractions, configuration formats, and execution assumptions [derouiche2025agenticaiframeworksarchitectures; kapoor2025ai]. As a result, organizations struggle when porting agents across stacks, evaluating behaviors consistently, or reusing designs developed by other teams. These challenges hinder reproducibility, slow down prototyping-to-deployment, and complicate governance in enterprise environments [10.1145/3711896.3736570; chen2025aiopslab].
Concurrently, several standardization efforts are shaping the agent ecosystem. Protocols such as Anthropic’s Model Context Protocol (MCP) [mcp] focus on resource and tool provisioning; inter-agent communication initiatives like Google’s Agent2Agent [a2a] and BeeAI’s Agent Communication Protocol (ACP) [acp] standardize message exchange. In addition, the Agntcy initiative [agntcy] aims to promote open interoperability and governance standards for the emerging agent ecosystem, providing shared principles for agent definition, discovery, and interaction across frameworks. While these efforts advance interoperability at the resource and communication layers, there still exists a critical gap at the level of agent behavior and execution semantics. In practice, this means there is no common standard for how control and data flows are managed, how agent state evolves, or how inputs and outputs are validated, leading to inconsistencies in agent execution behavior across different runtime environments.
To address these gaps, we introduce Agent Spec, a declarative, framework-agnostic configuration language for defining AI agents and their workflows with high fidelity. By providing a unified representation, Agent Spec enables seamless reusability, portability and interoperability across diverse frameworks, ensuring consistent behavior and reproducibility.
In practice, Agent Spec serves as an abstraction layer above framework-specific implementations, providing a unifying interface that captures agent functionality beyond individual framework constraints. It makes evaluation a first-class concern by defining common execution semantics and measurement protocols, so that agent and workflow performance can be compared across frameworks - similar in spirit to how HELM [liang2023holistic] and related harnesses standardized LLM evaluation. Like HELM, EleutherAI’s lm-evaluation-harness [eleutherai2024lmEvalHarness] and MT-Bench [bai2024mtbench101], Agent Spec centers evaluation as a core principle. However, while these efforts focus on model-centric standardized assessment, Agent Spec standardizes agent- and workflow-centric standardized assessment across frameworks. By fixing a common intermediate representation, execution semantics, and measurement protocol, Agent Spec supports reproducible, cross-platform comparisons of agent quality, safety, robustness, and efficiency, thus bridging the gap between agent portability, reusability and rigorous evaluation.
Our contributions in this work are as follows:
  • We present Agent Spec, a declarative, framework-agnostic language for specifying agents and workflows with high fidelity, enabling define-once, run-anywhere portability and seamless reusability across diverse runtimes.
  • We formalize a core set of components and execution semantics - Agents, Flows, Nodes, Tools, and control/data flow edges and discuss design principles for reliable, composable, and reusable agentic systems.
  • We release accompanying toolsets: a Python SDK (PyAgentSpec  [^1] https://pypi.org/project/pyagentspec/) for programmatic authoring and serialization, a reference runtime (WayFlow  [^2] https://github.com/oracle/wayflow) with native Agent Spec support, and runtime adapters for popular frameworks to demonstrate cross-framework execution.
  • Agent Spec serves as an Evaluation harness, enabling consistent comparison of agent and flow patterns across different frameworks. Using this capability, we report initial results on representative tasks from three benchmark datasets. Interestingly, we observe that both LangGraph and WayFlow deliver reliable accuracy and query time stability across datasets, AutoGen outperforms on complex reasoning tasks (e.g., on $\tau^{2}$-Bench [barres2025tau2]) but with increased latency, and CrewAI’s multi-agent coordination yields limited accuracy gains despite longer query times.
2 Background & Motivation
2.1 Related Work
AI agent frameworks provide the building blocks for constructing agentic systems that can reason, make decisions, and complete multi-step tasks. In recent years, with the advent of large language models (LLMs), several general-purpose frameworks have emerged with distinct abstractions, focus areas, and strengths, including LangChain [LangChain], LangGraph [LangGraph], GPT-Functions [openaiFunctionCalling], AutoGen [autogen], OCI Agents [ociagents], Semantic Kernel [semanticKernel], and CrewAI [crewAI]. These frameworks can augment LLMs (e.g., through Agents, Flows) with structured planning [wang2022self], memory [wang2023augmenting], tool use [schick2023toolformer], and the ability to perform search and access external environments [zhu2023large; kim2024rada].
LangChain [LangChain] offers a flexible, modular SDK for composing applications with LLMs, tools, long-term memory, and retrieval, and it underpins a large ecosystem of integrations. LangGraph [LangGraph] builds on LangChain to model agents and workflows as stateful, directed graphs with explicit control flow, enabling deterministic branching, loops, and multi-actor coordination; its associated tooling (e.g., LangGraph Studio) provides visualization and debugging support. AutoGen [autogen] focuses on multi-agent conversations with tool use and human-in-the-loop capabilities, providing abstractions for autonomous agent roles, inter-agent messaging, and extensible tool execution, which makes it suitable for complex, iterative workflows. It is designed with an open and flexible conversational communication style between agents, and its event-driven and distributed architecture makes it suitable for workflows that require long-running agents that collaborate across information boundaries with variable degrees of human involvement. Semantic Kernel [semanticKernel] is a model-agnostic SDK (with support across multiple programming languages) that provides planners, memories, connectors, and orchestration utilities to integrate LLMs and data sources into production applications. OCI AI Agent Platform [ociagents] targets enterprise-grade deployments with built-in tools and patterns (e.g., natural language SQL generation, retrieval-augmented generation), emphasizing scalability, manageability, and integration with enterprise services. CrewAI [crewAI] centers on multi-agent collaboration with explicit roles, tasks, and shared state, offering opinionated primitives for coordination and memory.
Separately, GPT-Functions (tool/function calling) [openaiFunctionCalling] specifies how models such as GPT-4o can invoke user-defined tools through structured function signatures, enabling safer and more controllable LLM-tool interaction across many runtimes. OctoTools [lu2025octotools] proposes an extensible framework oriented toward complex reasoning with “tool cards” that standardize wrappers around heterogeneous tools to improve transparency, reduce integration errors, and simplify maintenance.
While these frameworks advance agentic capabilities in complementary directions - ranging from SDKs to orchestration, coordination, and tool interoperability - the ecosystem remains fragmented. Differences in configuration models, component semantics, and execution assumptions hinder reusability, reliability, and reproducibility. This fragmentation motivates a unifying, declarative, and framework-agnostic specification such as Agent Spec, which provides a reusable & portable representation of agents and workflows executable across diverse runtimes.
2.2 Motivation
@@
Existing agentic frameworks have their own strengths and limitations (e.g., limited tool types, lack of generalization across diverse tasks) [raheem2025agentic; huang2025ai; joshi2025review]. While flexibility is crucial to accommodating diverse agentic design patterns, currently available frameworks do not fully support all of them. Table 1 outlines key objectives that a comprehensive framework should cover and describes how Agent Spec addresses each of these requirements.
@@
Table 1: Addressing challenges and objectives in agentic frameworks through the Open Agent Specification.
@@
Agents Author and edit conversational solutions such as ReAct [yao2023react] style agents. Define “resource” components (LLMs, memory, tools, or agents) attachable to other components’ properties.
Flows Author and edit graph-style workflows such as multi-step business processes. Define components with well-defined execution paths (such as directed graphs), supporting conditional or looping routes.
Control flow routing Ability to define the flow of execution between components in the system, including conditional or looping flows. Define control flow edges enabling one-to-many or many-to-one transitions where the specific transition to be taken is determined at runtime for a given execution.
I/O routing Ability to define how outputs feed inputs for fixed, conditional, or looping flows, including reuse across steps. Define data flow edges supporting one-to-many or many-to-one transitions where the specific transition to be taken is determined at runtime for a given execution.
Nesting & Composition Reuse components or solutions at any level of granularity in other solutions; includes multi-agent composition. Define components that can be encapsulated and reused; control message visibility, and orchestration of execution order and dependency either statically or dynamically.
Referencing Ability to define a component instance once and refer to it in multiple places. Use reference syntax to specify property values and reuse defined components across multiple flows or tools. For example, define an Agent once and reference it via an AgentNode in a Flow.
@@
Moreover, different frameworks are usually parameterized and configured by different means. This makes porting solutions between frameworks a tedious and error-prone process, hindering collaboration and knowledge sharing among teams developing agentic applications.
To address these challenges, a declarative, shareable, portable, and framework-independent representation of agentic design is essential. Analogous to how ONNX [onnx] standardized ML model interchange across frameworks such as PyTorch [pytorch] and TensorFlow [tensorflow], Agent Spec aims to define a common interchange format for AI agents and workflows. Just as ONNX enables models trained in one framework to be executed in another without reimplementation, Agent Spec allows agents to be authored once and deployed consistently across multiple frameworks.
2.3 Agent Spec within the Ecosystem
Agent Spec streamlines the design of agentic assistants and workflows by providing a unified, framework-agnostic configuration language that defines agent patterns, components, and behaviors. It complements, rather than replacing, ongoing standardization efforts such as:
  • Anthropic’s Model Context Protocol (MCP) [mcp]: Standardizes resource and data provision via client/server based RESTful APIs.
  • Google’s Agent2Agent Protocol [a2a] and BeeAI’s Agent Communication Protocol (ACP) [acp]: Standardized APIs for distributed agent communication.
  • The Agntcy initiative [agntcy] fosters open interoperability and governance standards, defining shared principles for agent discovery and cross-framework collaboration.
While protocols like MCP and A2A standardize tool or resource provisioning as well as inter-agent communication, Agent Spec complements these efforts by enabling standardized configuration of components related to agentic system design and execution in general, as shown in Fig. 1.
Figure 1: Position of Agent Spec within the emerging agent-standardization stack. MCP governs resource and data provisioning, A2A/ACP specify inter-agent communication, and Agent Spec defines the declarative layer for agent behavior and execution semantics.
3 Agent Spec: Objectives and Contributions
@@ 3.1 Objectives of Open Agent Specification
Agent Spec defines a common representation that generalizes the capabilities for AI agents and workflows. Agent Spec graphs can be executed in different frameworks through runtimes that map its components to framework-specific primitives. Combined with importers and exporters for various frameworks, this enables transparent portability of agentic workflows. Agent Spec also provides validation mechanisms to ensure that workflows are syntactically and logically correct before execution, promoting consistency and reliability.
3.2 Key Contributions of Agent Spec
Agent Spec offers several key contributions that address current challenges in the development and deployment of agentic workflows. The following subsections highlight how Agent Spec improves portability, reusability, interoperability, and unification across diverse frameworks, ultimately enhancing the effectiveness and robustness of AI agent systems.
@@ 3.2.1 Portability & Reusability
Framework-Agnostic: Agent Spec abstracts agent definitions from specific implementations by providing a unified declarative representation, allowing agents to operate uniformly across multiple platforms (e.g., AutoGen, LangGraph, WayFlow) without any need for reimplementation.
Modular Design: Its component-based architecture promotes reusability and extensibility, allowing complex agentic systems to be composed from standardized primitives.
Support for Complex Workflows: Agent Spec supports advanced workflows and modular design patterns, facilitating the seamless development of sophisticated AI agents.
Portable Execution Graphs: Agent Spec enables execution graphs to be described in a way that is independent of the underlying runtime implementation, ensuring seamless portability across frameworks. Converters between the representations of other frameworks to Agent Spec simplifies the porting of previously developed agents to Agent Spec-compatible frameworks. A recent example is deprecation of AutoGen framework and corresponding migration to MS Agent Framework [autogenmigration].
@@ 3.2.2 Interoperability & Compatibility
Robustness & Consistency: Agents defined in Agent Spec can be built on top of components that have been defined and validated within the Agent Spec representation, leading to higher reliability and predictability.
Define Once, Run on Multiple Frameworks: Agentic systems leveraging Agent Spec can run on various frameworks through runtime adapters, fostering open interoperability. This open specification allows developers to implement these adapters to support any framework, enabling agents to run in multiple environments.
@@ 3.2.3 Unification across Frameworks
Common Format: Agent Spec’s common interchange format simplifies integration of agents and enables easier agent sharing across platforms.
Knowledge Exchange: The unified declarative format that Agent Spec provides helps developers as well as maintainers of agents to efficiently exchange agent ideas and implementations.
Conformance Test Suite: The Agent Spec conformance test suite offers a comprehensive set of tests that cover the various functionalities and components available within the specification, such as Agent, Flow, BranchingNode, MapNode, and ToolNode. It includes illustrative validation examples of Agent Spec configurations and verifies that these configurations produce consistent execution outcomes and behaviors across multiple runtime environments. This approach ensures interoperability and correctness of implementations supporting the Agent Spec specification.
4 Design Principles and Key Components
This section provides an overview of the declarative model defined by Agent Spec. It explains how Agent Spec operates across multiple frameworks and outlines design details including SDK structure, supported programming languages, agent definitions using the SDK, and serialization and deserialization processes for agents and workflows. Figure 2 provides an overview of the Agent Spec ecosystem.
Agent Spec is a portable, shareable, platform-agnostic configuration language that describes agents and agentic systems with sufficient fidelity for execution across runtimes. It defines the conceptual objects, called components, that make up agents in typical agentic systems, together with the properties that configure their behavior and semantics.
Runtimes implement these Agent Spec components for execution within their own frameworks. Agent Spec is supported by SDKs in multiple languages (starting with Python), enabling serialization and deserialization of Agents to and from JSON (or YAML), or direct creation from object representations while maintaining full specification compliance. As shown in Table 2, Agent Spec provides for AI agents what ONNX provides for ML models.
Figure 2: Agent Spec’s Design.
4.1 Overview of Agent Specification
Table 2: Agent Spec serves a similar unifying role for AI agents as ONNX does for ML models.
@@ 4.1 Overview of Agent Specification
Scope Unified representation of ML models, allowing portability across different deep learning frameworks. Unified representation of AI agents and workflows, enabling them to run across diverse agent frameworks.
Portability ONNX allows ML models to be trained in one framework (e.g., PyTorch, TensorFlow) and executed in another. Agent Spec enables AI agents to be built in one framework (e.g., LangGraph [LangGraph], AutoGen [autogen], crewAI [crewAI]) and run in another one.
Unification Provides a common declarative format which is versioned as an Opset. Provides a common agent configuration format with versioning support.
Extensibility Supports various ML operations and optimizations. Defines modular components for scalable agent systems.
The Open Agent Specification defines structure and behavior of the conceptual building blocks, called components, that make up agents in typical agent-based systems. The following sections describe each component in greater detail; the full specification can be found in [language_spec_25_4_1].